# Dispersion of first sound in a weakly interacting ultracold Fermi liquid: an exact calculation

### v1

@@ Abstract
Introduction— Landau’s Fermi liquid theory is an effective theory which, when it is applicable, greatly simplifies the description of the many-body dynamics of a system of fermions, breaking it down to a single kinetic equation on a distribution of dilute quasiparticles. It is a very successful theory in describing the phenomenology of a wide class of fermionic systems, such as liquid 3He Baym and Pethick (1991); Vollhardt and Wolfle (1990), electron gases Jones and McClure (1966); Behnia (2022); Belitz and Kirkpatrick (2022), quantum gases Vichi (2000), down to nuclear/neutron matter Schäfer (2014). Nevertheless, once the phenomenological parameters characterizing the quasiparticle dispersion and interaction function are known, either from measurements Greywall (1986); Matsumoto et al. (1996) or from a microscopic calculation Lipschitz and Pitaevskii (1981a), Fermi liquid theory provides quantitative predictions on dynamical properties, such as the transport coefficients Højgård Jensen et al. (1968); Brooker and Sykes (1968).
@@
In normal 3He, despite decades of research, there is still a discrepancy between theory and experiment on the value of the transport coefficients. Consider for instance the shear viscosity $\eta$, which scales with temperature as $T^{-2}$: theory still underestimate the product $\eta T^{2}$ from the measurement by more than $20\%$ Abel et al. (1966); Baym and Pethick (1991); Parpia, J. M. et al. (1978); González et al. (2016); Baten et al. (2023). This is due to a limited knowledge of the quasiparticle interaction function, which has not be been computed from a microscopic theory, and whose experimental determination is limited to the lowest spherical harmonics Baym and Pethick (1991). As a consequence, the exact solutions of the transport equation Højgård Jensen et al. (1968); Brooker and Sykes (1968) were never validated experimentally, and relaxation time approximations Khalatnikov and Abrikosov (110); Baym and Pethick (1991) remain in use today González et al. (2016).
@@
Ultracold gases of fermionic gases provide exciting opportunities to quantitatively test those transport calculations Huang et al. (2024). These gases behave as Fermi liquids when the $s$-wave scattering length $a$ is negative and sufficiently small to open a regime of temperatures $T_{c}\ll T\ll T_{\textrm{F}}$ where $T_{c}$ is the superfluid critical temperature and $T_{\textrm{F}}$ the Fermi temperature. A Fermi liquid regime may also exist when $a$ is positive and small enough to suppress three-body recombination Ji et al. (2022). The quasiparticle dispersion and interaction function can be computed perturbatively in powers of $k_{\textrm{F}}|a|$ (with $k_{\textrm{F}}$ the Fermi wavenumber) Lipschitz and Pitaevskii (1981a). Experimentally, both the interaction strength and the temperature can be varied such that the typical collision time $\tau$ can be adjusted over several orders of magnitude Riedl et al. (2008); Patel et al. (2020), to explore both collisionless and hydrodynamic regimes Huang et al. (2024). Flat-bottom potentials Gaunt et al. (2013); Navon et al. (2021), where sound can be excited at very low wavevector $q$ in homogeneous samples allow to study the propagation and attenuation of sound waves in a very controlled environment Patel et al. (2020).
@@
Theoretically, great efforts were devoted to the calculation of the viscosity at strong coupling, in particular in the unitary regime $|a|=+\infty$ Rupak and Schäfer (2007); Bruun and Smith (2007); Enss et al. (2011) and exact results are available in the high temperature virial regime Braby et al. (2010); Nishida (2019); Fujii and Enss (2023). At intermediate temperatures however, a controlled approach has not been found due to the absence of a separation of timescales between the collisional and kinetic dynamics Lipschitz and Pitaevskii (1981b). The temperature range of the Fermi liquid regime shrinks as the quasiparticle cross section increases with the interaction strength Pavaskar and Rothstein (2022), and it is eventually hidden by the onset of a superfluid phase at a critical temperature $T_{c}\approx 0.17T_{\textrm{F}}$ at unitarity. Below $T_{c}$, sound attenuation is dominated by phonon-phonon interactions Rupak and Schäfer (2007); Enss et al. (2011); Kurkjian et al. (2017). In the weakly interacting normal phase, the transport coefficients were computed using relaxation time approximations Vichi (2000); Massignan et al. (2005); Watabe et al. (2010); Maki and Enss (2023), (either in the “variational” Massignan et al. (2005); Maki and Enss (2023), or in the original Abrikosov-Khalatnikov Vichi (2000); Watabe et al. (2010) formulation), even though their uncontrolled error increases towards low temperatures. In this Letter, we perform an exact calculation of the transport coefficients Højgård Jensen et al. (1968); Brooker and Sykes (1968) to lowest order in $k_{\textrm{F}}|a|$, and show that the error of the relaxation time approximations is significant, up to 25%.
@@
For negative values of $a$, the quasiparticle interactions are attractive Recati and Stringari (2010), which prevents the emergence of a zero sound mode as in liquid 3He. We thus lack a parameter similar to $c_{0}-c_{1}$, where $c_{0}$, $c_{1}$ are the speed of zero and first sound respectively, to characterize the dispersion of sound, as was done in 3He Abel et al. (1966). In this work, we derive the leading-order deviation of the frequency $\omega_{q}$ from its linear depart $c_{1}q$. To do so, we solve the transport equation to the order $\tau^{2}$ of the Chapman-Enskog’s expansion. This is the so-called second hydrodynamic order often used in a relativistic context to cure the acausality of the diffusion equations characteristic of dissipative hydrodynamics Israel (1976). We find the exact solution of the transport equation by decomposing the quasiparticle distribution function on a basis of orthogonal polynomials adapted to the low temperature limit Hofmann and Gran (2023). Remarkably, the frequency shift involves only two parameters of the collision kernel: the viscous relaxation time $\tau_{\eta}$ and a second-order viscous time $t_{\eta}$, which we both compute exactly. As for the dispersion of the sound branch Kurkjian et al. (2017), we find that $\omega_{q}$ is above its linear depart $c_{1}q$, the deviation being proportional to $q^{3}\tau^{2}$ with $\tau$ scaling as $1/a^{2}T^{2}$.
@@
where $\delta n_{\sigma}=n_{\sigma}-n^{\rm eq}_{\sigma}$ is the fluctuation about thermal equilibrium $n_{\sigma}^{\rm eq}(\epsilon_{\rm eq})=1/(1+\textrm{e}^{(\epsilon_{\rm eq}-\mu)/T})$. In the general case, the equilibrium dispersion relation $\epsilon_{\rm eq}$ and interaction function $f_{\sigma\sigma^{\prime}}$ are phenomenological parameters usually rexpressed in terms of an effective mass and Landau parameters. In the case of a weakly interacting gas with contact interactions, these quantities can be calculated perturbatively in powers of the coupling constant $g=4\pi a/m$ Lipschitz and Pitaevskii (1981a) (we use $\hbar=k_{\rm B}=1$ throughout this work). To first order in perturbation theory, we have:
@@
For $T\ll T_{\rm F}$, transport occurs in a energy shell of typical depth $T$ around the Fermi energy $\epsilon_{\rm F}$ Nozières and Pines (1966); Baym and Pethick (1991). We thus reparametrize the quasiparticle distribution as
@@
We choose $Q_{0}=1$ and $Q_{1}=\epsilon$ as the initial condition. Note that even and odd polynomials are respectively symmetric and antisymmetric about the Fermi surface: $Q_{n}(-\epsilon)=(-1)^{n}Q_{n}(\epsilon)$. By contrast to this exact approach, the relaxation time approximations Khalatnikov and Abrikosov (110); Baym and Pethick (1991) truncate the expansion in Eq. (13) to $n=0$, thereby neglecting the energy dependence of the quasiparticle distribution. The difference between the Abrikosov-Khalatnikov Khalatnikov and Abrikosov (110) and the variational formulation Baym and Pethick (1991); Massignan et al. (2005); Braby et al. (2010) lies in the treatment of the remaining $\epsilon$ dependence of the collision kernel: Abrikosov and Khalatnikov replaced it by its value in $\epsilon=0$ (in particular they approximate the quasiparticle lifetime by its value at the Fermi level $\Gamma(\epsilon)\approx\Gamma(0)$), while the variational formulation averages it over $\epsilon$.
Note that the polynomials $Q_{n}$ used here differ from the orthogonal polynomials of the momentum $p$ used at nonvanishing temperature Braby et al. (2010); Huang et al. (2024), and that the replacement Eq. (6) converts [^1] Denoting $H_{n}$ and $T_{n}$ the polynomials defined respectively by the first and second line of Eq. (S85) of the supplementary material of Ref. Huang et al. (2024) (so as to avoid the confusion with the $Q_{n}$ polynomials defined in this work), one has $Q_{n}(\epsilon)=\lim_{m\to+\infty}H_{2n}(\sqrt{\epsilon+m})=\lim_{m\to+\infty}\frac{H_{2n+1}(\sqrt{\epsilon+m})}{\sqrt{\epsilon+m}}=\lim_{m\to+\infty}\frac{T_{2n+2}(\sqrt{\epsilon+m})}{m}$ even/odd powers of $p^{2}$ into even/odd powers of $\epsilon$.
@@
Note that the subspaces of symmetric and antisymmetric functions of $\epsilon$ (even and odd $n$ respectively) are decoupled: $\Gamma_{nn^{\prime}}=\mathcal{N}_{nn^{\prime}}^{l}=0$ if ${n+n^{\prime}}$ is odd Brooker and Sykes (1970). This is specific to the low-temperature limit where both the density of state $g(\epsilon)$ and energy integration domain are symmetric about the Fermi surface. Treating separately the odd and even orders, $\Gamma$ and $\mathcal{S}$ are tridiagonal matrices and can be expressed analytically as
@@
$\mathcal{S}_{nn^{\prime}}\!\!\!$ $=$ $\!\!\!\left[2\pi^{2}\frac{n^{2}+n-1}{4n^{2}+4n-3}\right]\delta_{nn^{\prime}}+\frac{\delta_{n-2,n^{\prime}}}{n(n-1)}$ (19)
$\qquad\qquad\qquad\qquad+\frac{\delta_{n+2,n^{\prime}}\xi_{n+2}\xi_{n+1}}{(n+2)(n+1)}$
@@
where we have introduced the matrix $\mathcal{M}^{2}=(\mathcal{M}_{nn^{\prime}}^{l=2})_{nn^{\prime}}$ and the unit vectors $(\vec{u}_{n})_{n^{\prime}}=\delta_{nn^{\prime}}$. We have neglected in Eq. (26) the vector $\vec{\nu}^{3}$ and the components of $\vec{\nu}^{1}$ orthogonal to the zero eigenvector $\nu_{0}^{1}\vec{u}_{0}$: both are of order $O(\omega_{0}\tau)^{2}$ (see Fig. 1), and hence negligible compared to the subleading term $c\vec{\nu}^{2}$ in Eq. (26). As is noted in Ref. Brooker and Sykes (1970), the expansion in powers of $\omega_{0}\tau$ becomes very tedious as soon as higher Legendre components have to be taken into account. Fortunately this is not the case in our calculation to order $O(\omega_{0}\tau)^{2}$: the transport equation can be truncated to $l=2$, and only two collision times are needed to express $\nu_{0}^{2}$ from Eq. (26):
@@
where we introduce the viscous collision time $\tau_{\eta}/\tau=\vec{u}_{0}\frac{1}{\mathcal{M}^{2}}\vec{u}_{0}$ and the second-order viscous time $t_{\eta}/\tau=\sqrt{\vec{u}_{0}\frac{1}{(\mathcal{M}^{2})^{2}}\vec{u}_{0}}$. These two parameters characterize the exact equations of motion of the density and parallel velocity within second-order hydrodynamics; a relaxation time approximation would not distinguish them since it amounts to replacing $\mathcal{M}^{2}$ by a number. In an exact calculation, they can be expressed as infinite series Brooker and Sykes (1968); Højgård Jensen et al. (1968) or as continued fractions Hofmann and Das Sarma (2022), which converge rather slowly when truncated at $n=n^{\prime}=n_{\rm max}$: $\tau_{\eta}-\tau_{\eta}^{(n_{\rm max})}=O(1/n_{\rm max}^{2})$. Numerically, we obtain
@@
Figure 2: The sound attenuation and dispersion coefficients $\gamma$ and $\Delta c$ obtained by fitting a Lorentzian function $c\mapsto Z/\left[(c-\bar{c}_{1}-(\omega_{0}\tau)^{2}B(\omega_{0}\tau))^{2}+(\omega_{0}\tau)^{2}A(\omega_{0}\tau)^{2}\right]$ (green and orange lines in inset, respectively for $\omega_{0}\tau=0.04$ and $\omega_{0}\tau=2$) to the density response $\nu_{0}^{0}$ obtained by numerically solving Eq. (20) at non vanishing $\omega_{0}\tau$. The horizontal dashed and dotted line show respectively the exact hydrodynamic limits $B\to-\textrm{Im}(z_{1})/(\omega_{0}\tau)$ and $A\to\textrm{Re}(z_{1}-\bar{c}_{1})/(\omega_{0}\tau)^{2}$ from Eq. (30), and the relaxation time approximation ($\tau_{\eta},t_{\eta}\to\tau_{\eta}^{(0)}$). This figure is drawn in the limit $k_{\textrm{F}}a\to 0$ ($c_{1}/v_{\textrm{F}}=1/\sqrt{3}$).
@@
This exact value of $\eta$ recovers the result of Refs. Brooker and Sykes (1968); Højgård Jensen et al. (1968); Baym and Pethick (1991) for the case of an isotropic scattering amplitude (see Eq. (11) in Højgård Jensen et al. (1968) with $\alpha=2/5$). This is a factor $\tau_{\eta}/{\tau_{\eta}^{(0)}}\simeq 1.08$ above the value in the relaxation time approximation of Refs. Massignan et al. (2005); Bruun and Smith (2007); Braby et al. (2010) (called the “variational approximation” therein). We note that the underestimation consecutive to the approximation is larger than stated before Bruun and Smith (2007); Braby et al. (2010). Conversely, in the Abrikosov-Khalatnikov approximation Khalatnikov and Abrikosov (110); Nozières and Pines (1966); Watabe et al. (2010), the viscosity is overestimated by a factor ${\tau_{\eta}^{\rm(AK)}}/{\tau_{\eta}}=4\tau_{\eta}^{(0)}/3{\tau_{\eta}}\simeq 1.24$.
Experimentally, the attenuation and dispersion coefficients appearing in Eqs. (31)–(32) can be measured through the shape of the response functions. Fig. 2 shows that they can be accurately recovered from Lorentzian fits of the density response. The shift $\omega_{q}-c_{1}q$ remains positive throughout the crossover from the hydrodynamic to the collisionless regime, as visible on Fig. 3.
Figure 3: The transition from hydrodynamic (red curves) to collisionless regime (blue curves) in the velocity response function $\nu_{n=0}^{l=1}$ of a weakly interacting Fermi gas with $k_{\textrm{F}}a=-0.5$. The broadening and the positive shift of the resonance frequency from the first sound velocity (vertical dashed line, Eq. (25)) is clearly visible at small $\omega_{0}\tau$. The curve are obtained by numerically solving Eq. (20) truncated to $n_{\rm max}=100$ and $l_{\rm max}$ ranging from 20 in the hydrodynamic regime to 8500 in the collisionless regime.
@@
Finally, we note that the nonvanishing isotropic components of the collision kernel $\mathcal{M}^{0,\perp}=(\mathcal{M}^{0})_{n>1,n^{\prime}>1}$, which are usually associated with the bulk viscosity $\zeta$, do not affect the propagation of sound, even within second-order hydrodynamics. The decoupling is a consequence of the parabolic dispersion relation (2) of the Fermi liquid quasiparticles, since this does not couple the total velocity $\nu_{0}^{1}$ directly to the nonconserved isotropic components of the quasiparticle distribution $\nu_{n>1}^{0}$. This means the bulk viscosity appears in the hydrodynamic equations only at higher temperatures and stronger interactions. Nevertheless, the comparison of the exact to approximate matrix elements $(1/\mathcal{M}^{0,\perp})_{22}\simeq 1.41/\mathcal{M}^{0}_{22}$ suggests that the error made by the relaxation time approximations on the value of the bulk viscosity is large.

### v2

@@ Abstract
Introduction— Landau’s Fermi liquid theory is an effective theory which, when it is applicable, greatly simplifies the description of the many-body dynamics of a system of fermions, breaking it down to a single kinetic equation on a distribution of dilute quasiparticles. It is a very successful theory in describing the phenomenology of a wide class of fermionic systems, such as liquid 3He [1, 2], electron gases [3, 4, 5], quantum gases [6], down to nuclear/neutron matter [7, 44]. Nevertheless, if both the the quasiparticle interaction function and the collision probability [8] are known, either from measurements [9, 10] or from a microscopic calculation [11], Fermi liquid theory provides quantitative predictions on dynamical properties, such as the transport coefficients [12, 13, 14].
@@
In normal 3He, despite decades of research, there is still a discrepancy between theory and experiment on the value of the transport coefficients. Consider for instance the shear viscosity $\eta$, which scales with temperature as $T^{-2}$: theory still underestimate the product $\eta T^{2}$ from the measurement by more than $20\%$ [15, 1, 16, 17, 18]. This is due to a limited knowledge of the quasiparticle interaction function and collision probability [1], which were not computed from a microscopic theory, and whose experimental determination is limited to the lowest spherical harmonics. As a consequence, the exact solutions of the transport equation [12, 13, 14] were never validated experimentally, and relaxation time approximations [19, 1] remain in use today [17].
@@
Ultracold gases of fermionic gases provide exciting opportunities to quantitatively test those transport calculations [20]. These gases behave as Fermi liquids when the $s$-wave scattering length $a$ is negative and sufficiently small to open a regime of temperatures $T_{c}\ll T\ll T_{\textrm{F}}$ where $T_{c}$ is the superfluid critical temperature and $T_{\textrm{F}}$ the Fermi temperature. A Fermi liquid regime may also exist when $a$ is positive and small enough to suppress three-body recombination [21]. The quasiparticle dispersion and interaction function can be computed perturbatively in powers of $k_{\textrm{F}}|a|$ (with $k_{\textrm{F}}$ the Fermi wavenumber) [11]. Experimentally, both the interaction strength and the temperature can be varied such that the typical collision time $\tau$ can be adjusted over several orders of magnitude [22, 23], to explore both collisionless and hydrodynamic regimes [20]. Flat-bottom potentials [24, 25], where sound can be excited at very low wavevector $q$ in homogeneous samples allow to study the propagation and attenuation of sound waves in a very controlled environment [23].
@@
Theoretically, great efforts were devoted to the calculation of the viscosity at strong coupling, in particular in the unitary regime $|a|=+\infty$ [26, 27, 28] and exact results are available in the high temperature virial regime [29, 30, 31]. At intermediate temperatures however, a controlled approach has not been found due to the absence of a separation of timescales between the collisional and kinetic dynamics [32]. The temperature range of the Fermi liquid regime shrinks as the quasiparticle cross section increases with the interaction strength [33], and it is eventually hidden by the onset of a superfluid phase at a critical temperature $T_{c}\approx 0.17T_{\textrm{F}}$ at unitarity. Below $T_{c}$, sound attenuation is dominated by phonon-phonon interactions [26, 28, 34]. In the weakly interacting normal phase, the transport coefficients were computed using relaxation time approximations [6, 35, 36, 37], (either in the “variational” [35, 37], or in the original Abrikosov-Khalatnikov [6, 36] formulation), even though their uncontrolled error increases towards low temperatures. In this Letter, we perform an exact calculation of the transport coefficients [12, 13] to lowest order in $k_{\textrm{F}}|a|$, and show that the error of the relaxation time approximations is significant, up to 25%.
@@
For negative values of $a$, the quasiparticle interactions are attractive [38], which prevents the emergence of a zero sound mode as in liquid 3He. We thus lack a parameter similar to $c_{0}-c_{1}$, where $c_{0}$, $c_{1}$ are the speed of zero and first sound respectively, to characterize the dispersion of sound, as was done in 3He [15]. In this work, we derive the leading-order deviation of the frequency $\omega_{q}$ from its linear depart $c_{1}q$. To do so, we solve the transport equation to the order $\tau^{2}$ of the Chapman-Enskog’s expansion. This is the so-called second hydrodynamic order often used in a relativistic context to cure the acausality of the diffusion equations characteristic of dissipative hydrodynamics [39]. We find the exact solution of the transport equation by decomposing the quasiparticle distribution function on a basis of orthogonal polynomials adapted to the low temperature limit [40]. Remarkably, the frequency shift involves only two parameters of the collision kernel: the viscous relaxation time $\tau_{\eta}$ and a second-order viscous time $t_{\eta}$, which we both compute exactly. As for the dispersion of the sound branch [34], we find that $\omega_{q}$ is above its linear depart $c_{1}q$, the deviation being proportional to $q^{3}\tau^{2}$ with $\tau$ scaling as $1/a^{2}T^{2}$.
@@
where $\delta n_{\sigma}=n_{\sigma}-n^{\rm eq}_{\sigma}$ is the fluctuation about thermal equilibrium $n_{\sigma}^{\rm eq}(\epsilon_{\rm eq})=1/(1+\textrm{e}^{(\epsilon_{\rm eq}-\mu)/T})$. In the general case, the equilibrium dispersion relation $\epsilon_{\rm eq}$ and interaction function $f_{\sigma\sigma^{\prime}}$ are phenomenological parameters usually rexpressed in terms of an effective mass and Landau parameters. In the case of a weakly interacting gas with contact interactions, these quantities can be calculated perturbatively in powers of the coupling constant $g=4\pi a/m$ [11] (we use $\hbar=k_{\rm B}=1$ throughout this work). To first order in perturbation theory, we have:
@@
For $T\ll T_{\rm F}$, transport occurs in a energy shell of typical depth $T$ around the Fermi energy $\epsilon_{\rm F}$ [41, 1]. We thus reparametrize the quasiparticle distribution as
@@
We choose $Q_{0}=1$ and $Q_{1}=\epsilon$ as the initial condition. Note that even and odd polynomials are respectively symmetric and antisymmetric about the Fermi surface: $Q_{n}(-\epsilon)=(-1)^{n}Q_{n}(\epsilon)$. By contrast to this exact approach, the relaxation time approximations [19, 1] truncate the expansion in Eq. (13) to $n=0$, thereby neglecting the energy dependence of the quasiparticle distribution. The difference between the Abrikosov-Khalatnikov [19] and the variational formulation [1, 35, 29] lies in the treatment of the remaining $\epsilon$ dependence of the collision kernel: Abrikosov and Khalatnikov replaced it by its value in $\epsilon=0$ (in particular they approximate the quasiparticle lifetime by its value at the Fermi level $\Gamma(\epsilon)\approx\Gamma(0)$), while the variational formulation averages it over $\epsilon$.
Note that the polynomials $Q_{n}$ used here differ from the orthogonal polynomials of the momentum $p$ used at nonvanishing temperature [29, 20], and that the replacement Eq. (6) converts even/odd powers of $p^{2}$ into even/odd powers of $\epsilon$.
@@
Note that the subspaces of symmetric and antisymmetric functions of $\epsilon$ (even and odd $n$ respectively) are decoupled: $\Gamma_{nn^{\prime}}=\mathcal{N}_{nn^{\prime}}^{l}=0$ if ${n+n^{\prime}}$ is odd [42]. This is specific to the low-temperature limit where both the density of state $g(\epsilon)$ and energy integration domain are symmetric about the Fermi surface. Treating separately the odd and even orders, $\Gamma$ and $\mathcal{S}$ are tridiagonal matrices and can be expressed analytically as
@@
$\mathcal{S}_{nn^{\prime}}\!\!\!$ $=$ $\!\!\!{2\pi^{2}\frac{n^{2}+n-1}{4n^{2}+4n-3}}\delta_{nn^{\prime}}\!+\!\frac{\delta_{n-2,n^{\prime}}}{n(n-1)}\!+\!\frac{\delta_{n+2,n^{\prime}}\xi_{n+2}\xi_{n+1}}{(n+2)(n+1)}$
@@
where we have introduced the matrix $\mathcal{M}^{2}=(\mathcal{M}_{nn^{\prime}}^{l=2})_{nn^{\prime}}$ and the unit vectors $(\vec{u}_{n})_{n^{\prime}}=\delta_{nn^{\prime}}$. We have neglected in Eq. (25) the vector $\vec{\nu}^{3}$ and the components of $\vec{\nu}^{1}$ orthogonal to the zero eigenvector $\nu_{0}^{1}\vec{u}_{0}$: both are of order $O(\omega_{0}\tau)^{2}$ (see Fig. 1), and hence negligible compared to the subleading term $c\vec{\nu}^{2}$ in Eq. (25). As is noted in Ref. [42], the expansion in powers of $\omega_{0}\tau$ becomes very tedious as soon as higher Legendre components have to be taken into account. Fortunately this is not the case in our calculation to order $O(\omega_{0}\tau)^{2}$: the transport equation can be truncated to $l=2$, and only two collision times are needed to express $\nu_{0}^{2}$ from Eq. (25):
@@
where we introduce the viscous collision time $\tau_{\eta}/\tau=\vec{u}_{0}\frac{1}{\mathcal{M}^{2}}\vec{u}_{0}$ and the second-order viscous time $t_{\eta}/\tau=\sqrt{\vec{u}_{0}\frac{1}{(\mathcal{M}^{2})^{2}}\vec{u}_{0}}$. These two parameters characterize the exact equations of motion of the density and parallel velocity within second-order hydrodynamics; a relaxation time approximation would not distinguish them since it amounts to replacing $\mathcal{M}^{2}$ by a number. In an exact calculation, they can be expressed as infinite series [13, 12] or as continued fractions [43], which converge rather slowly when truncated at $n=n^{\prime}=n_{\rm max}$: $\tau_{\eta}-\tau_{\eta}^{(n_{\rm max})}=O(1/n_{\rm max}^{2})$. Numerically, we obtain
@@
Our value coincides with the exact calculations of $\eta$ done in the context of 3He [14] or neutron matter [44]. This is a factor $\tau_{\eta}/{\tau_{\eta}^{(0)}}\simeq 1.08$ above the value in the relaxation time approximation of Refs. [35, 27, 29] (called the “variational approximation” therein). We note that the underestimation consecutive to the approximation is larger than stated before [27, 29]. Conversely, in the Abrikosov-Khalatnikov approximation [19, 41, 36], the viscosity is overestimated by a factor ${\tau_{\eta}^{\rm(AK)}}/{\tau_{\eta}}=4\tau_{\eta}^{(0)}/3{\tau_{\eta}}\simeq 1.24$.
@@
Finally, we note that the nonvanishing isotropic components of the collision kernel $\mathcal{M}^{0,\perp}=(\mathcal{M}^{0})_{n>1,n^{\prime}>1}$, which are usually associated with the bulk viscosity $\zeta$, do not affect the propagation of sound, even within second-order hydrodynamics. This means the bulk viscosity appears in the hydrodynamic equations of a Fermi liquid. Nevertheless, the comparison of the exact to approximate matrix elements $(1/\mathcal{M}^{0,\perp})_{22}\simeq 1.41/\mathcal{M}^{0}_{22}$ suggests that the relaxation time approximations estimate the bulk viscosity quite poorly.